18. In a redox titration, potassium permanganate, KMnO4, was standardized using a 19.60 g/L

FeSO4(NH4)SO4.6H2O, (molar mass = 392.21 g/mol) standard solution. A 20.00 mL of the

standard solution, acidified with H2SO4 was put in a conical flask. A titre volume 15.50

mL of KMnO4 solution was required to reach the end-point. The titration reaction is:

Fe2+(aq) + MnO4

−(aq) → Fe3+(aq) + Mn2+(aq)

(i) Balance the above redox reaction in basic medium.

(ii) Calculate the molarity of the standard solution.

(iii) Determine molarity of MnO4-(aq) in the burette.


1
Expert's answer
2022-04-18T17:04:09-0400

(i) Balance the above redox reaction in acidic medium. (Basic medium is a mistake, the solution is acidified with H2SO4).


Fe2+(aq) + MnO4−(aq) → Fe3+(aq) + Mn2+(aq)


Fe2+(aq) - e- → Fe3+(aq) x5

MnO4(aq) + 8H+ + 5e- → Mn2+(aq) + 4H2O x1


5Fe2+(aq) + MnO4−(aq) + 8H+ → 5Fe3+(aq) + Mn2+(aq) + 4H2O


(ii) Calculate the molarity of the standard solution.


c(Fe2+) = 19.60/392.21 = 0.05 M


(iii) Determine molarity of MnO4-(aq) in the burette.


20 mL = 0.020 L

n(Fe2+) = 0.05*0.020 = 0.001 mol => n(KMnO4) = n(Fe2+)/5 = 0.0002 mol

15.50 mL = 0.01550 L

c(KMnO4) = 0.0002/0.01550 =0.0129 M